OpenStudy (anonymous):

the price of shares in a company is falling by an average of 2% per day. what percentage of their value are there after 2 days? (geometrical applications)

OpenStudy (tkhunny):

Unclear. Falling by an average of 2% over what time period? Annually?

OpenStudy (tkhunny):

Okay, 2% per day for 2 days. What say you?

OpenStudy (texaschic101):

It fell 4%.....right ?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

If it was 200 when it started, where would it be after one day?

OpenStudy (texaschic101):

How do you know it was 200 when it started ?

OpenStudy (tkhunny):

I don't, it's just a question. Answer, please. You have 30 seconds.

OpenStudy (texaschic101):

after one day it would be (.02)(200) = 4 200 - 4 = 196 right ?

OpenStudy (dwelsher):

on the second day what is .02 of 98?

OpenStudy (tkhunny):

Excellent. Now do the EXACT same thing with 196. Do NOT skip the first step because you think you already know what it is. Do it ALL.

OpenStudy (dwelsher):

What do you mean? do the same thing with 196. .02 of 196=?

OpenStudy (tkhunny):

Correct. EXACT same thing. Or, just answer the question, if it was 196 on one day, what will it be on the next day?

OpenStudy (dwelsher):

196-(.02*196)=? i'm not supposed to be giving answers to every problem

OpenStudy (texaschic101):

(.02)(196) = 3.92 196 - 3.92 = 192.06

OpenStudy (dwelsher):

right on!

OpenStudy (texaschic101):

I still don't understand where you were getting the 200 from ?

OpenStudy (dwelsher):

I assume from a question

OpenStudy (texaschic101):

read the original question.....

OpenStudy (dwelsher):

I see what you mean. I assumed on the first day you had 100% and subtracted 2%.then you went to 98 and figured .02 % of that

OpenStudy (texaschic101):

wouldn't you have to know the original price of the shares to determine what percentage of their value is left after 2 days ?

OpenStudy (texaschic101):

does it start with 100% ?

OpenStudy (tkhunny):

Okay, now stop doing it in two steps. One Day in Two Steps. 200 * 0.02 = 4 200 - 4 = 196 One Day in One Step 200 - (200*0.02) = 200(1 - 0.02) = 200*0.98 Two Days in One Step \(200*(0.98)^{2}\) Okay, that's enough background and playing around with my made-up numbers. Answer the question. You do NOT need the original price. I made up 200 just to have something concrete. It matters not. Answer the question.

OpenStudy (dwelsher):

so after 2 days you have 96.34% of what ever value. Does this make sense.?

OpenStudy (texaschic101):

(.02)(100) = 2 100 - 2 = 98 (.02)(98) = 1.96 98 - 1.96 = 96.04%

OpenStudy (texaschic101):

What am I doing wrong.....I didn't get what you got ?

OpenStudy (dwelsher):

yes you are right- good job

OpenStudy (texaschic101):

thanks....I understand now.....I feel stupid for not catching on sooner

OpenStudy (tkhunny):

Where is @yasmin95 ?

OpenStudy (texaschic101):

I don't know....she isn't answering

OpenStudy (dwelsher):

I doubt if a texaschic 101 with a smart score of 70 would have much problems witha question like this

OpenStudy (texaschic101):

I just didn't know what to start with.....it now makes sense to start with 100%

OpenStudy (dwelsher):

it would have come to you